OpenStudy (anonymous):

What is the solution of the equation? 10 + sqrt x + 8 = -4 A. 44 B. 28 C. -2 D. 36

OpenStudy (anonymous):

\[10+\sqrt{(x+8)}=-4\]or \[10+\sqrt{x}+8=-4\]

OpenStudy (anonymous):

the second one @KinzaN

OpenStudy (anonymous):

Then you begin by subtracting 10 and then 8 from both sides: \[10 -10 +\sqrt{x} +8 -8 = -4 - 10 - 8\]

OpenStudy (anonymous):

Okay, I don't get the -10 and -8 under the -4.. Do I have to subtract both from the -4?

OpenStudy (anonymous):

@KinzaN

OpenStudy (anonymous):

If your question looks like the second one I posted, then you have to subtract 10 and 8 from both sides to "move" the numbers over to isolate for "x"

OpenStudy (anonymous):

Haha, oh okay.

OpenStudy (anonymous):

That makes more sense. \[-10+\sqrt{x+8}=-4\]

OpenStudy (anonymous):

Sorry, I just get confused by all the parenthesis and all.

OpenStudy (anonymous):

So you have to add 10 to both sides as a first step, following the rules of BEDMAS (but backwards) to isolate for "x" that's under the root.

OpenStudy (anonymous):

I have to add 10 to 10 and 8 or the 10 and -4? and what's BEDMAS? lol

OpenStudy (anonymous):

\[-10+(10)+\sqrt{x+8}=-4 +(10)\]

OpenStudy (anonymous):

Well you're isolating for x, right? So the order of operations you're following are: BEDMAS... Brackets Exponents Division Multiplication Addition Subtraction but backwards!--so "SAMDEB"

OpenStudy (anonymous):

Meaning you do your addition/subtraction first, then multiplication/division, followed by exponents and then brackets.

OpenStudy (anonymous):

so that would be sqrt x + 8 = 6

OpenStudy (anonymous):

\[(\sqrt{x+8})^{2} = 6^{2}\] \[x+8-(8)=36-(8)\] \[x=28\]

OpenStudy (anonymous):

Yeah! :D

OpenStudy (anonymous):

Oh!! Thanks! How did you get the ^2? was it because of the parenthesis? :)